OpenStudy (anonymous):

HELP! Determine Whether 4x^2-10 Is A Difference Of Two Squares. If So, Choose The Correct Factorization?.

OpenStudy (anonymous):

seems unlikely since 10 is not a perfect square

OpenStudy (anonymous):

then again since \(\sqrt{10}^2=10\) you could write \[(2x-\sqrt{10})(2x+\sqrt{10})\]

OpenStudy (anonymous):

Thanks Your A BIG Help ((: